trying to calculate how many liquid ounces will be contained in an item that is 1.5 inches wide, 1 inch tall and 11 inches long. The topic bar was not visible for me to make a selection.
Email address is Bonniecsu@yahoo.com
Thank You,
Bonnie

Answer by Alan3354(69443)

trying to calculate how many liquid ounces will be contained in an item that is 1.5 inches wide, 1 inch tall and 11 inches long. The topic bar was not visible for me to make a selection.
-------------
The volume = 1.5*1*11 = 16.5 cubic inches
I think 1 fl oz (a very stupid unit) = 29.58 cc
1 cubic inch = 2.54^3 cc = 16.387 cc
16.5 cubic in = 270.39 cc
270.39/29.58 = 9.14 fl oz
------------------
And people have problems understanding metric units